OpenStudy (addictwithapen21):

You roll a number cube twice. Find the probability of the event in fraction form. Need help with #2 and #3 1. Rolling a 4 then an even number. 1/12 2. Rolling a 3 then a 3 3. Rolling a number less than 1 and then a number less than 2

satellite73 (satellite73):

first one is right

satellite73 (satellite73):

second one, same idea. multiply

satellite73 (satellite73):

first one you got by\[\frac{1}{6}\times \frac{1}{2}\] right?

OpenStudy (addictwithapen21):

yeah

OpenStudy (sshayer):

there is no number <1

satellite73 (satellite73):

oh right

OpenStudy (sshayer):

in a cube die numbers are from 1,2,...,6

satellite73 (satellite73):

so i guess you are supposed to say "zero"

OpenStudy (addictwithapen21):

so #3 would be 0?

satellite73 (satellite73):

yes, impossible event has probability zero

OpenStudy (addictwithapen21):

okay and what about #2? how do i find the fractions i need to multiply

OpenStudy (agent0smith):

What's the probability of rolling a 3 when you roll a die?

OpenStudy (agent0smith):

#1 is harder than number 2, if you could do 1 you should be able to do 2.